A child releases a 6.25L helium balloon in the parking lot of an amusement park where the temperature is 28.50oC and the air pressure is 757.2mmHg, what will be the volume of the balloon when it has risen to an altitude where the temperature is -34.35oC and the air pressure is 366.4mmHg?
1
Expert's answer
2020-09-22T03:36:54-0400

The combined gas law states that the volume of a given amount of a gas is directly proportional to it's pressure and absolute temperature.


"\dfrac{P1V1}{T1}=\dfrac{P2V2}{T2}"


V1=6.25 L

T1=28+273=301 K

P1=757.2 mmHg or 0.996atm

V2=??

T2=-34.35+273=238.65 K

P2=366.4 mmHg or 0.482atm


Plug in the values into the formula;

"\dfrac{6.25*0.996}{301}=\dfrac{V2*0.482}{238.65}"


"0.021=0.002V2"


"V2=10.5L"